At 9:00 a.m. a truck leaves the truck yard and travels west at a rate of 50 mph. At 11:00 a.m., a second truck leaves along the same route, traveling at 75 mph. When will the second truck catch up to the first?
Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(31518)

At 9:00 a.m. a truck leaves the truck yard and travels west at a rate of 50 mph. At 11:00 a.m., a second truck leaves along the same route, traveling at 75 mph. When will the second truck catch up to the first?
------------------------
The 1st truck is 100 miles away (50 x 2)
The 2nd truck gains on the 1st at 25 mph (75- 50)
100/25 = 4 hours
11:00 + 4 = 1500 hours

Answer by rfer(12730)

A has 100 mi lead
over take speed is 25 mph
100/25=4 hrs
11:00 + 4= 3:00 pm